If any object is travelling at a constant acceleration i.e. 5m/s/s then the distance covered by that object after 4s as per the formula S=ut+0.5att is 40m but if I calculate the same through v=u+at the same is coming to 50m.


Will apprecite if you can explain where my understanding is wrong.

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Swarnendu Sengupta,

Ans:- I think you want a long walk in fresh air. It would refresh your mind and you will not do silly mistakes again.

S=0.5×5×4²=40 m   OK

Then v=0+(5×4)=20m/s

Let the ditance travelled by it is=X

v²=u²+2fX

20²=0+(2×5)X

X=40m

Hence X=S UNDERSTOOD ?

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Swarnendu Sengupta !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur